I have a solution of 3.1875 meters for this problem but it is in error. (I calculated height/distance from lamppost to the man as 8/15; calculated height/distance from man to end of shadow as 1.7/x). I appreciate your help!
A man is walking away from a lamppost with a light source 8 meters above the ground. The man is 1.7 meters tall. How long is the man's shadow when he is 15 meters from the lamppost?
Thanks, Linda
: I have a solution of 3.1875 meters for this problem but it is in error. (I calculated height/distance from lamppost to the man as 8/15; calculated height/distance from man to end of shadow as 1.7/x). I appreciate your help!
A man is walking away from a lamppost with a light source 8 meters above the ground. The man is 1.7 meters tall. How long is the man's shadow when he is 15 meters from the lamppost?
Thanks, Linda

Answer by oscargut(667)

(8/(15+x))=(1.7/x)
then
8x=1.7(15+x)
8x=25.5+1.7x
6.3x=25.5
x=25.5/6.3=4.047619
Answer: 4.0477 m (aprox)
Question 155516: I have a solution of 3.1875 meters for this problem but it is in error. (I calculated height/distance from lamppost to the man as 8/15; calculated height/distance from man to end of shadow as 1.7/x). I appreciate your help!
A man is walking away from a lamppost with a light source 8 meters above the ground. The man is 1.7 meters tall. How long is the man's shadow when he is 15 meters from the lamppost?
Thanks, Linda
: I have a solution of 3.1875 meters for this problem but it is in error. (I calculated height/distance from lamppost to the man as 8/15; calculated height/distance from man to end of shadow as 1.7/x). I appreciate your help!
A man is walking away from a lamppost with a light source 8 meters above the ground. The man is 1.7 meters tall. How long is the man's shadow when he is 15 meters from the lamppost?
Thanks, Linda

Answer by stanbon(18983)

A man is walking away from a lamppost with a light source 8 meters above the ground. The man is 1.7 meters tall. How long is the man's shadow when he is 15 meters from the lamppost?
------------------------
Draw the picture.
The lamppost is the height of the larger triangle;
The man is height of the smaller SIMILAR triangle.
----------------
Let the base of the larger triangle be x.
-------------------
EQUATION:
8 meters/ 1.7 meters = x/x-15
8(x-15) = 1.7x
8x - 120 = 1.7x
6.3x = 120
x = 19.048
----------------
Man's shadow = 19.048-15 = 4.048 meters
=========================================
Cheers,
Stan H.
Question 155516: I have a solution of 3.1875 meters for this problem but it is in error. (I calculated height/distance from lamppost to the man as 8/15; calculated height/distance from man to end of shadow as 1.7/x). I appreciate your help!
A man is walking away from a lamppost with a light source 8 meters above the ground. The man is 1.7 meters tall. How long is the man's shadow when he is 15 meters from the lamppost?
Thanks, Linda
: I have a solution of 3.1875 meters for this problem but it is in error. (I calculated height/distance from lamppost to the man as 8/15; calculated height/distance from man to end of shadow as 1.7/x). I appreciate your help!
A man is walking away from a lamppost with a light source 8 meters above the ground. The man is 1.7 meters tall. How long is the man's shadow when he is 15 meters from the lamppost?
Thanks, Linda

Answer by Earlsdon(3748)

You can use the priniples of similar triangles here.
Let the length of the man's shadow be x meters.
The 8-meter lampost is the height of the first right triangle while the 1.7-meter man is the height of the second right triangle.
The base of the first triangle is 15+x meters while the base of the second right triangle is x meters, which is what we are trying to find.
The rule is "Corresponding sides of similar triangles are proportional"
So we can write the proportion:
8/(x+15) = 1.7/x Do you see this? Now we solve for x, the length of the man's shadow by cross-multiplying.
8x = 1.7(x+15) Simplifying, we get:
8x = 1.7x+25.5 Subtract 1.7x from both sides.
6.3x = 25.5 Divide both sides by 6.3
x = 4.047
The length of the man's shadow is 4.047 meters (approximately).